OpenStudy (anonymous):

is the square root of -64 = -8?

OpenStudy (anonymous):

\[\sqrt{-64^{?}}\] = -8 ?

OpenStudy (anonymous):

\[\sqrt{-64} = \sqrt{64} \cdot \sqrt{-1}\]\[\sqrt{-1} = i\]

OpenStudy (anonymous):

so its the answer -8i?

OpenStudy (anonymous):

Just 8i

OpenStudy (anonymous):

oh is it because of the (sqrt)-1 ?

OpenStudy (anonymous):

Yes. We factor the negative one out to get i. Then we take the square root of a positive number.

OpenStudy (anonymous):

great tytytytyty :)